1. For each of the following annuities calculate the annual cash flow

Present Value

Years

Interest Rate

$24,500 6 11%
$19,700 8 7%

2.  You deposit $5,000 at the end of each year for 20 years into an accouninterest, how much money will you have in the account in 20 years?

3. Lycan Inc. has 7% coupon bonds on the market that have 9 years left to maturity. The bonds make annual payments and have a par value of $1,000. If the YTM on these bonds is 8.4%, what is the current bond price?

4. The Timberlake Co. has 7% coupon bonds in the market with 9 years left to maturity. The bonds make annual payments and have a par value of $1,000. If the bonds currently sell for $961.50, what is the YTM?

5. Volbeat Co. has bonds on the market with 10.5 years to maturity, a YTM of 6.2%, a par value of $1,000, and a current price of $945. The bonds make semiannual payments.

What must the coupon rate be on the bonds?
Answer #1

Present value =   24500
  
Years (n) =   6
  
Interest rate (i)= 11% or    0.11
  
Equal annuity payments formula

= P* i *((1+i)^n)/(((1+i)^n)-1)  
=24500*0.11*((1+0.11)^6)/(((1+0.11)^6)-1)  
=5791.225808  
So, equal annual cash flow is    $5,791.23

(2)

Present value =   19700
  
Years (n) =   8
  
Interest rate (i)= 7% or    0.07
  
Equal annuity payments formula = P* i *((1+i)^n)/(((1+i)^n)-1)  
19700*0.07*((1+0.07)^8)/(((1+0.07)^8)-1)  
3299.114921  
So, equal annual cash flow is    $3,299.11

(One question is answered as per HomeworkLib policy)